If 2^(2x) = 7 find x.
Found 2 solutions by MathLover1, Alan3354:
Answer by MathLover1(20850)
Answer by Alan3354(69443)

If 2^(2x) = 7 find x.
--------
I found it. It's after the 2 inside the parentheses.
---
Ok, seriously:
2x*log(2) = log(7)
x = log(7)/(2log(2))) = log(7)/log(4)
========================
You've posted several problems that are similar.
I didn't notice the others until after I had done this one.
You should learn how to solve them, not just get the answers.